# Symmetry and Topology of Successive Quantum Feedback Control

### v1

@@ Abstract
Meanwhile, quantum feedback control has attracted intensive interest from several perspectives. Continuous measurement and feedback [20] has long been central, as it enables state stabilization [21, 22, 23, 24, 25], noise suppression [26, 27, 28, 29], feedback cooling [30, 31, 24, 32, 33], and error correction [34, 35, 36, 37, 38, 39]. Moreover, quantum feedback control has intrinsic connections to quantum thermodynamics as illustrated by Maxwell’s demon, which clarifies the role of information in thermodynamics both theoretically [40, 41, 42, 43, 33] and experimentally [44, 45, 46, 47, 48, 32]. However, the topological aspects of quantum feedback control have been rarely studied. A first step toward such a framework for topological quantum feedback control has been presented in Ref. [19], while the results of symmetry classification therein are limited to the case where only a single projective measurement without error is executable.
In this Letter, we provide a general symmetry classification of successive feedback control. We prove a no-go theorem stating that, despite the greatly larger protocol space compared to the single projective measurement case, the admissible BL classes again collapse to the AZ classes. Our central assumption is that each measurement is bare [49], meaning that with positive Kraus operators. As a consequence, some symmetries that might appear available in generic non-Hermitian evolutions are ruled out. We illustrate the topology of a chiral Maxwell’s demon subject to Gaussian measurement errors, for which we observe quantized winding numbers. Since any POVM admits a realization via a bare measurement, our result broadly applies and provides design principles for robust quantum feedback protocols, based on topology and symmetry.
@@
Setup.— We consider quantum feedback control, composed of measurement and feedback. In general, the measurement is described by Kraus operators $\{\hat{M}_{m}\}$ satisfying $\sum_{m}\hat{M}_{m}^{\dagger}\hat{M}_{m}=\hat{I}$, and the feedback is described by unitary operators conditioned on the outcomes of preceding measurements. In our setup (see also Fig. 1), the CPTP map for the entire process is written as
@@
where $\left\{\hat{M}_{m_{i}}^{(i)}\right\}$ is the set of measurement operators at step $i$, and $\hat{U}_{m_{1},\cdots,m_{i}}^{(i)}(\tau_{i})$ is the feedback unitary conditioned on outcomes $m_{1},\cdots,m_{i}$ with feedback time duration $\tau_{i}$.
In general, the POVM $\{\hat{E}_{m}\}$ associated with the Kraus operators $\{\hat{M}_{m}\}$ is given by $\hat{E}_{m}=\hat{M}_{m}^{\dagger}\hat{M}_{m}$. Then, each Kraus operator has the polar decomposition $\hat{M}_{m}=\hat{V}_{m}\sqrt{\hat{E}_{m}}$ with $\hat{V}_{m}$ being some unitary. There is a certain arbitrariness whether this unitary $\hat{V}_{m}$ is regarded as a part of measurement or a part of feedback. Here, we suppose that the feedback in our setup does not include such unitary but is only given by the continuous-time evolution generated by a Hamiltonian, written as
@@
In the main theorem (stated later) of this work, we further assume that the unitary $\hat{V}_{m}$ is absent in the measurement and thus the Kraus operator is simply given by $\hat{M}_{m}=\sqrt{\hat{E}_{m}}$ ($\geq 0$). We refer to this class of measurements as bare measurements [49]. We note that $\hat{M}_{m}=\sqrt{\hat{E}_{m}}$ is a unique choice to make the Kraus operator positive. It is known that such class can be characterized as measurements with minimal disturbance [20, 50].
In addition, we assume that the sequence of measurements are non-adaptive throughout this paper. That is, whereas feedback is dependent on the preceding measurement outcomes, the measurement itself is fixed in advance and does not depend on preceding outcomes.
@@
We now consider a more detailed setup relevant to topological classifications. Consider a single particle on a $d$-dimensional lattice with translational symmetry. Under periodic boundary conditions (PBC), $\mathcal{E}$ can be block-diagonalized in the eigenspaces of translation operators. Each block corresponds to the Bloch matrix [19] defined by
@@
On the other hand, we assumed that every symmetry sector has an eigenvalue $1$. This spectral structure forbids certain BL symmetries. Specifically, if $\mathcal{E}$ has an eigenvalue $\xi$ and has either $P$ symmetry or $C$ symmetry with $\epsilon_{C}=-1$, $\xi$ is also an eigenvalue. Likewise, if $\mathcal{E}$ has $K$ or $Q$ symmetry, $\epsilon_{K}\xi^{*}$ or $\epsilon_{Q}\xi^{*}$ is also an eigenvalue. Consequently, successive bare measurements are incompatible with $P$ symmetry or any $\chi$ symmetry with $\epsilon_{\chi}=-1$ ($\chi=C,K,Q$). Taken together with some intrinsic constraints on the BL symmetries and the assumption that $\mathcal{E}$ has no residual unitary symmetry, this result yields precisely the ten-fold AZ classes in Table 1 [19].
@@
Figure 2: Spectra of the quantum channels of chiral Maxwell’s demon with projective measurement (left) and that with Gaussian errors (right). The colored dots are the eigenvalues under PBC for size $L=400$, and the color corresponds to the wavenumber $k$ of the Bloch matrices. The unit of energy, the height of the potential and the feedback duration are set to be $J=1,V=10^{2}$, $\tau=1$ respectively. We set the radius of the error (cutoff) to be $c=2$, and the Gaussian sharpness is set to be $\ell=15$. The winding number $w$ around the origin is $w(\xi_{\text{PG}}=0)=-1$ for the left and $w(\xi_{\text{PG}}=0)=-2$ for the right, and that around $1-\epsilon$ is $w(\xi_{\text{PG}}=1-\epsilon)=-1$ for both pictures with $\epsilon=10^{-12}$. The highly degenerate eigenvalue $\xi=0$ is omitted in both figures.
@@
Figure 2 shows the eigenspectrum of the CPTP map $\mathcal{E}$ of chiral Maxwell’s demon with Gaussian errors. The result is for finite system size $L$, but we observe the robustness of the spectrum against increase in $L$ (see End Matter), which justifies the extrapolation to $L\to\infty$. We can observe that the high degeneracy at $\xi=0$ is partially broken. This is because the error of the measurement keeps off-diagonal elements $|\outerproduct{i}{j}$ with $|i-j|\leq 2c$ nonzero, leaving $4c+1$ bands. Notably, these additional bands around the origin change the winding number around $\xi_{\text{PG}}=0$. By contrast, the spectral structure around $\xi=1$ is insensitive to the measurement error, which motivates us to alternatively choose the base-point at $\xi_{\text{PG}}=1-\epsilon$ with sufficiently small $\epsilon>0$. The reason why we need $\epsilon$ is that the spectrum of any CPTP map exactly passes $1$ due to the trace-preserving condition, while the base-point should be avoided by the spectrum when defining the winding number. Another motivation for $\xi_{\text{PG}}=1-\epsilon$ is the analogy with topology of one-dimensional classical stochastic processes [15, 17, 16], where the natural choice of the base-point corresponds to the steady state.
@@
We remark that the extra bands near the origin have a quantum origin; some off-diagonal degrees of freedom survive by the measurement error and contribute to the new bands. This cannot happen with classical stochastic dynamics because it does not make off-diagonal elements alive, leaving only one band in the complex plane. This sensitivity to off-diagonal elements may be exploited to elucidate the role of quantum coherence in quantum feedback control.
@@
We also note that this model can be applied to continuous measurement regime by taking $\ell$ proportional to $\tau$ and letting $\tau\to 0$ [55]. However, in this perspective it is no longer valid to appropriately take the cutoff $c$ of the Gaussian tail, and therefore, the winding number might not be well-defined since the size of the Bloch matrices diverges as $L\to\infty$, accompanied by an infinite number of bands scattered on the complex plane. Nevertheless, from the observed behavior of the spectrum against $c$, we expect that those infinite bands are condensed around 0. If the only accumulation point of the spectrum is 0, we are still able to define the winding number around $\xi_{\text{PG}}=1-\epsilon$.
@@
$\hat{\sigma}_{z}X_{\text{trunc}}(k)\hat{\sigma}_{z}=-X_{\mathrm{trunc}}(k).$ (18)
@@
We conclude this Letter with some future prospects. First, it remains to be discussed whether the AZ restriction holds when one allows adaptive measurements, where one can decide what measurement to perform depending on the earlier outcomes. Moreover, considering that some measurement processes such as photon counting inevitably possess unitary component followed by the bare part [20], it is essential to discuss what happens beyond the bare measurement assumption. It is still unclear whether the CPTP condition alone casts some intrinsic constraint on the classification. Here, we conjecture that all the 38 BL symmetry classes are accessible by engineering the unitary parts of non-bare measurements.

### v2

@@ Abstract
Meanwhile, quantum feedback control has attracted intensive interest from several perspectives. Continuous measurement and feedback [20] has long been central, as it enables state stabilization [21, 22, 23, 24, 25], noise suppression [26, 27, 28, 29], feedback cooling [30, 31, 24, 32, 33], and error correction [34, 35, 36, 37, 38, 39]. Moreover, quantum feedback control has intrinsic connections to quantum thermodynamics as illustrated by Maxwell’s demon, which clarifies the role of information in thermodynamics both theoretically [40, 41, 42, 43, 33] and experimentally [44, 45, 46, 47, 48, 32]. However, the topological aspects of quantum feedback control have rarely been studied. A first step toward such a framework for topological quantum feedback control has been presented in Ref. [19], while the results of symmetry classification therein are limited to the case where only a single projective measurement without error is executable.
In this Letter, we provide a general symmetry classification of successive feedback control. We prove a no-go theorem stating that, despite the greatly larger protocol space compared to the single projective measurement case, the admissible BL classes again collapse to the AZ classes. Our central assumption is that each measurement is bare [49], meaning that all Kraus operators are positive. As a consequence, some symmetries that might appear available in generic non-Hermitian evolutions are ruled out. We illustrate the topology of a chiral Maxwell’s demon subject to Gaussian measurement errors, for which we observe quantized winding numbers. Since any POVM admits a realization via a bare measurement, our result broadly applies and provides design principles for robust quantum feedback protocols based on topology and symmetry.
@@
Setup.— We consider quantum feedback control composed of measurement and feedback. In general, the measurement is described by Kraus operators $\{\hat{M}_{m}\}$ satisfying $\sum_{m}\hat{M}_{m}^{\dagger}\hat{M}_{m}=\hat{I}$, and the feedback is described by unitary operators conditioned on the outcomes of preceding measurements. In our setup (see also Fig. 1), the CPTP map for the entire process is written as
@@
where $\left\{\hat{M}_{m_{i}}^{(i)}\right\}$ is the set of measurement operators at step $i$, and $\hat{U}_{m_{1},\cdots,m_{i}}^{(i)}(\tau_{i})$ is the feedback unitary conditioned on outcomes $m_{1},\cdots,m_{i}$ with feedback duration $\tau_{i}$.
In general, the POVM $\{\hat{E}_{m}\}$ associated with the Kraus operators $\{\hat{M}_{m}\}$ is given by $\hat{E}_{m}=\hat{M}_{m}^{\dagger}\hat{M}_{m}$. Then, each Kraus operator has the polar decomposition $\hat{M}_{m}=\hat{V}_{m}\sqrt{\hat{E}_{m}}$ with $\hat{V}_{m}$ being some unitary. There is a certain arbitrariness regarding whether this unitary $\hat{V}_{m}$ is regarded as a part of measurement or a part of feedback. Here, we suppose that the feedback in our setup does not include such unitaries, but is only given by the continuous-time evolution generated by a Hamiltonian, written as
@@
In the main theorem (stated later) of this work, we further assume that the unitary $\hat{V}_{m}$ is absent in the measurement and thus the Kraus operator is simply given by $\hat{M}_{m}=\sqrt{\hat{E}_{m}}$ ($\geq 0$). We refer to this class of measurements as bare measurements [49]. We note that $\hat{M}_{m}=\sqrt{\hat{E}_{m}}$ is a unique choice making the Kraus operator positive. It is known that such a class can be characterized as measurements with minimal disturbance [20, 50].
In addition, we assume that the sequence of measurements is non-adaptive throughout this paper. That is, whereas feedback is dependent on the preceding measurement outcomes, the measurement itself is fixed in advance and does not depend on preceding outcomes.
@@
We now consider a more detailed setup relevant to topological classifications. Consider a single particle on a $d$-dimensional lattice with translational symmetry. Under periodic boundary conditions (PBC), $\mathcal{E}$ can be block-diagonalized in the eigenspaces of the translation operators. Each block corresponds to the Bloch matrix [19] defined by
@@
On the other hand, we assumed that every symmetry sector has the eigenvalue $\xi=1$. This spectral structure forbids certain BL symmetries. Specifically, if $\mathcal{E}$ has an eigenvalue $\xi$ and has either $P$ symmetry or $C$ symmetry with $\epsilon_{C}=-1$, $\xi$ is also an eigenvalue. Likewise, if $\mathcal{E}$ has $K$ or $Q$ symmetry, $\epsilon_{K}\xi^{*}$ or $\epsilon_{Q}\xi^{*}$ is also an eigenvalue. Consequently, successive bare measurements are incompatible with $P$ symmetry or any $\chi$ symmetry with $\epsilon_{\chi}=-1$ ($\chi=C,K,Q$). Taken together with some intrinsic constraints on the BL symmetries and the assumption that $\mathcal{E}$ has no residual unitary symmetry, this result yields precisely the ten-fold AZ classes in Table 1 [19].
@@
Figure 2: Spectra of the CPTP maps of chiral Maxwell’s demon with projective measurement (a) and that with Gaussian errors (b.1) – (b.4). The colored dots are the eigenvalues under PBC for size $L=400$, whose colors correspond to the wavenumber $k$ of the Bloch matrices, and the black dots are the eigenvalues under OBC for size $L=20$. The unit of energy, the height of the potential and the feedback duration are set to be $J=1,V=10^{2}$, $\tau=1$ respectively. We set the radius of the error (cutoff) to be $c=2$, and the Gaussian sharpness is set to be $\ell=15$. The winding number $w$ around the origin is $w(\xi_{\text{PG}}=0)=-1$ for the left and $w(\xi_{\text{PG}}=0)=0$ for the right, and that around $1-\epsilon$ is $w(\xi_{\text{PG}}=1-\epsilon)=-1$ for both pictures with $\epsilon=10^{-12}$.
@@
Figure 2 shows the eigenspectrum of the CPTP map $\mathcal{E}$ of chiral Maxwell’s demon with Gaussian errors. The result is for finite system size $L$, but we observe the robustness of the spectrum against an increase in $L$ (see End Matter), which justifies the extrapolation to $L\to\infty$. We can observe that the high degeneracy at $\xi=0$ is partially broken. This is because the error of the measurement keeps the off-diagonal elements $|\outerproduct{i}{j}$ with $|i-j|\leq 2c$ nonzero, leaving $4c+1$ bands. Notably, these additional bands around the origin change the winding number around $\xi_{\text{PG}}=0$. By contrast, the spectral structure around $\xi=1$ is insensitive to the measurement error, which motivates us to alternatively choose the base-point at $\xi_{\text{PG}}=1-\epsilon$ with sufficiently small $\epsilon>0$. The reason why we need $\epsilon$ is that the spectrum of any CPTP map exactly passes $1$ due to the trace-preserving condition, while the base-point should be avoided by the spectrum when defining the winding number. Another motivation for $\xi_{\text{PG}}=1-\epsilon$ is the analogy with the topology of one-dimensional classical stochastic processes [15, 17, 16], where the natural choice of the base-point corresponds to the steady state.
@@
We remark that the extra bands near the origin have a quantum origin; some off-diagonal degrees of freedom survive due to measurement error and contribute to the new bands. This cannot happen with classical stochastic dynamics because it does not make off-diagonal elements alive, leaving only one band in the complex plane. This sensitivity to off-diagonal elements may be exploited to elucidate the role of quantum coherence in quantum feedback control.
As shown by black dots in Fig. 2, the loop structure of the spectrum under PBC collapses under the open boundary condition (OBC) for both the cases with and without errors. This implies that the non-Hermitian skin effect [1, 4, 56, 57] of quantum feedback control [19] persists under measurement errors.
@@
We also note that this model can be applied to the continuous measurement regime by taking $\ell$ proportional to $\tau$ and letting $\tau\to 0$ [55]. However, from this perspective it is no longer valid to appropriately take the cutoff $c$ of the Gaussian tail, and therefore, the winding number might not be well-defined since the size of the Bloch matrices diverges as $L\to\infty$, accompanied by an infinite number of bands scattered on the complex plane. Nevertheless, based on the observed behavior of the spectrum against $c$, we expect that those infinite bands are condensed around 0. If the only accumulation point of the spectrum is 0, we are still able to define the winding number around $\xi_{\text{PG}}=1-\epsilon$.
@@
$\sigma_{z}X_{\text{trunc}}(k)\sigma_{z}=-X_{\mathrm{trunc}}(k).$ (18)
@@
We conclude this Letter with some future prospects. First, it remains to be discussed whether the AZ restriction holds when one allows adaptive measurements, where one can decide what measurement to perform depending on the earlier outcomes. Moreover, considering that some measurement processes, such as photon counting, inevitably possess a unitary component followed by the bare part [20], it is essential to discuss what happens beyond the bare measurement assumption. It is still unclear whether the CPTP condition alone casts some intrinsic constraint on the classification. Here, we conjecture that all the 38 BL symmetry classes are accessible by engineering the unitary parts of non-bare measurements.